Construct a sampling distribution of sample mean and answer the questions on
your answer sheet.
Samples of 3 cards are drawn from a population of five cards numbered from 1-5.
1. How many are the possible outcomes?
2. What are the possible means?
3. What is the probability of getting 4 as a mean?
4. What is the probability of getting 2 as a mean?
5. What is the probability of getting 3.33 as a mean?
1
Expert's answer
2021-08-10T12:02:42-0400

N=5

n=3

Without replacement

1. Number of outcomes "= \frac{N!}{n!(N-n)!}"

"= \frac{5!}{3!(5-3)!} = \frac{4 \times 5}{2} = 10"

2.



The sampling distribution of sample mean



3. P(Mean=4) "= \frac{1}{10}=0.1"

4. P(Mean = 2) "= \frac{1}{10}=0.1"

5. P(Mean=3.33) "= \frac{2}{10}=0.2"